OpenStudy (anonymous):

3/8w - 22= -9/8w +4 ki dont understand this

hero (hero):

Lolz, someone posted this two seconds ago.

OpenStudy (anonymous):

ya that is my sister

hero (hero):

Place like terms on the same side.

OpenStudy (anonymous):

ok then what do you do after that

hero (hero):

9/8w + 3/8w = 22 + 4

hero (hero):

You know how to add fractions so don't tell me you don't know what to do next

OpenStudy (anonymous):

actually i dont know how to add fractions

hero (hero):

If you don't know how to add fractions, then you shouldn't be working on this type of problem.

OpenStudy (anonymous):

can you please help me i dont understand last year i feel asleep in class and this is my first year doing homeschool so know i have to learn this...

hero (hero):

To add fractions, you have to first make sure the denominators are the same. In this case, they're already the same, so we can combine them: (9+3)/8 w = 26 12w/8 = 26 12w = 8*26 4(3)w = 4(2*26) 3w = 52 w = 52/3

OpenStudy (anonymous):

kk thank you

hero (hero):

No, I'm not helping you with another one.

OpenStudy (anonymous):

can you also help me with this to thank you how do you simplify this 72/5